Jim started running track after school to get ready for basketball tryouts. The first day, he ran the quarter-mile track six times in 15 minutes. What was his rate of speed?
Answer by Alan3354(69443)

Jim started running track after school to get ready for basketball tryouts. The first day, he ran the quarter-mile track six times in 15 minutes. What was his rate of speed?
-------------
Rate of speed?
It's either his rate, or his speed.
-----
d = 6 times 1/4 mile = 1.5 miles
-----
d = r*t
r = d/t = 1.5 mi/15 min
r = 0.1 mi/minute